OpenStudy (anonymous):

Can someone tell me if I did this correctly? Essay: Show all work. A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 8y ft. and the middle length side x + 6y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simplest correct mathematical form, including units. x-8yft = shortest side L x+6yft= middle side length (x-8y)(x+6y)= (x)(x)=x^2 (x)(6y)=6xy (-8y)(x)=-8xy (-8y)(6y)=-48y^2 A=(x-8y)(x+6y)sq.ft. A=x^2-8xy+6xy-48yy^2 A=x^2-2xy-48y^2 sq. ft

jimthompson5910 (jim_thompson5910):

This would be true if it were a rectangle

jimthompson5910 (jim_thompson5910):

But it's a triangle. So you have to take half of this.

OpenStudy (anonymous):

how exaxtly do I take half of this? DO I just show my answer divided by two?

jimthompson5910 (jim_thompson5910):

You just take the final answer and divide by 2 like so \[\Large A = \frac{x^2-2xy-48y^2}{2}\]

jimthompson5910 (jim_thompson5910):

make sure to show the "divided by 2" part in every previous step to keep it consistent

OpenStudy (anonymous):

okay, so my answer would be to divide this out?

jimthompson5910 (jim_thompson5910):

yes, you can optionally keep going to get \[\Large A = \frac{x^2-2xy-48y^2}{2}\] \[\Large A = \frac{x^2}{2}-\frac{2xy}{2}-\frac{48y^2}{2}\] \[\Large A = \frac{1}{2}x^2-xy-24y^2\]

OpenStudy (anonymous):

ok, that makes sense:) ty

jimthompson5910 (jim_thompson5910):

yw